OpenStudy (anonymous):

please help. 1. sin x (csc(x)-sin (x)) 2. (1-cot^2(x))(1-cos^2(x)) 3. 1-tan^2(x) / sec^2(x) I have to simplify them but I'm not sure how

OpenStudy (anonymous):

So csc(x) is the same as 1/sin(x) cot(x) is the same as 1/tan(x) sec(x) is the same as 1/cos(x) tan(x) is the same as sin(x)/cos(x) Now that that is out of the way, there is a rule that sin^2(x) + cos^2(x) = 1 This equation can be changed to cos^2(x) = 1 - sin^2(x) or sin^2(x) = 1 - cos^2(x) You can also get two other equations out of this by dividing both sides by by sin^2(x) and cos^2(x) By dividing both sides by cos^2(x) you get sin^2(x)/cos^2(x) + 1 = 1/cos^2(x) This can be simplified to tan^2(x) +1 = sec^2(x) By dividing both sides by sin^2(x) you get 1 + cos^2(x)/sin^2(x) = 1/sin^2(x) This can be simplified to 1 + cot^2(x) = csc^2(x) I'm too lazy to do the rest but hopefully this helped with your problem.

OpenStudy (anonymous):

yeah it did. thanks a lot (: